OpenStudy (waheguru):

is s^1 the same as s

OpenStudy (wolfboy):

Yea because one s^1=s so they are the same

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no defference since any number to the power of 1 is a number itself.

OpenStudy (waheguru):

I was just making sure Thanks

OpenStudy (waheguru):

I have one more question

OpenStudy (waheguru):

Is s*s^4=s^5

OpenStudy (wolfboy):

Ummmm yea I think so cause s^5 is s*s*s*s*s so s*s^4 I think is the same